OpenStudy (anonymous):

Solving x^2+8x+6 by factoring?

OpenStudy (anonymous):

can not be done

OpenStudy (anonymous):

good luck with that program you are going to need the quadratic formula or what is easier in this case is completing the square

OpenStudy (anonymous):

Okay, then, how would I go about completing the square in this one?

OpenStudy (anonymous):

\[x^2+8x+6=0\] subtract \(6\) \[x^2+8x=-6\] then half of 8 is 4 so \[(x+3)^2=-6+16\\ (x+3)^2=10\]

OpenStudy (anonymous):

damn typo \[(x+4)^2=10\] is what i meant, ignore the 3

OpenStudy (anonymous):

that makes \[x=-4\pm\sqrt{10}\]

OpenStudy (anonymous):

Okay, no, I didn't understand that at all. But I don't know if that would have helped my objective anyway. I need to find the x and y intercepts.

OpenStudy (anonymous):

I'd love to just plug it into my calculator, but I have to do it algebraically.

OpenStudy (anonymous):

finding the x intercetps is the same as solving \[x^2+6x+8=0\] the \(y\) intercept is\((0,8)\) because if \(x=0\) then \(y=8\)

OpenStudy (anonymous):

Yeahh, but what are the x intercepts and how do I figure those out?